# Self-Tuning: Instructing LLMs to Effectively Acquire New Knowledge through Self-Teaching

### v2

@@ Abstract
Large language models (LLMs) often struggle to provide up-to-date information due to their one-time training and the constantly evolving nature of the world. To keep LLMs current, existing approaches typically involve continued pre-training on new documents. However, they frequently face difficulties in extracting stored knowledge. Motivated by the remarkable success of the Feynman Technique in efficient human learning, we introduce Self-Tuning, a learning framework aimed at improving an LLM’s ability to effectively acquire new knowledge from raw documents through self-teaching. Specifically, we develop a Self-Teaching strategy that augments the documents with a set of knowledge-intensive tasks created in a self-supervised manner, focusing on three crucial aspects: memorization, comprehension, and self-reflection. Additionally, we introduce three Wiki-Newpages-2023-QA datasets to facilitate an in-depth analysis of an LLM’s knowledge acquisition ability concerning memorization, extraction, and reasoning. Extensive experimental results on Llama2 family models reveal that Self-Tuning consistently exhibits superior performance across all knowledge acquisition tasks and excels in preserving previous knowledge.††∗Work done during the internship at Tencent AI Lab. 111Our source code and the curated new datasets can be accessed publicly at https://github.com/zhangxy-2019/Effective-Knowledge-Injection.
@@ 1 Introduction
In this paper, we focus on injecting new knowledge into the parameters of LLMs. As depicted in the upper part of Figure 1, a standard approach involves continued pre-training (A) on a raw corpus containing new information Jang et al. (2022). However, it struggles to extract the embedded knowledge, potentially due to the impaired question-answering (QA) capability Allen-Zhu and Li (2023); Cheng et al. (2024). Despite the assistance of subsequent instruction-tuning (B) Wei et al. (2022); Ouyang et al. (2022a) on QA data, the knowledge retrieved from the LLMs remains notably constrained Jiang et al. (2024b). Recently, Jiang et al. (2024b) suggests fine-tuning on a mix of QA data and related documents before continuing pre-training, with the aim of teaching the model how to access knowledge from documents and answer questions. Although this method greatly outperforms standard approaches, our initial results suggest that its effectiveness in knowledge extraction remains limited.
@@
We introduce Wiki-Newpages-2023-QA datasets to enable a comprehensive analysis of an LLM’s knowledge acquisition ability in terms of memorization, extraction, and reasoning.
@@ 4.1 Overview
Our objective is to equip an LLM with the ability to learn how to derive knowledge from raw documents. This is achieved by training the model using a combination of training document dataset and associated training QA dataset , as depicted in the upper left part of Figure 2. To enhance effective knowledge absorption, we present along with a series of knowledge-intensive tasks (a.k.a. self-teaching tasks) that are related to their content for Self-Teaching (in the lower part of Figure 2). These tasks are generated in a self-supervised manner based on the contents of , using the proposed Self-Teaching learning approach (Section 4.2). The multi-task training objective is presented as follows.
@@
Our aim is to train the model to apply the learned strategy for spontaneously extracting knowledge from unseen documents (i.e., the test document dataset ). In addition to training on , we include , allowing the model to review and refine its question-answering ability. This approach enhances knowledge extraction from . The training objective is as follows:
@@
Our goal is to empower the model to continuously assimilate new knowledge from previously unseen documents by conducting follow-up training on the dataset. The objective for this stage is as follows:
@@ 5.3 Training Dynamics
We analyze the training dynamics during continued pre-training on the test documents by altering the number of training epochs to rule out the possibility that the exceptional performance of Self-Tuning in enhancing knowledge acquisition is merely a result of early fitting on the test documents, and to provide a comprehensive examination of its long-term knowledge retention capability. Furthermore, we incorporate the results of PIT and continued pre-training for a holistic evaluation.
@@ 5.5 Variants of Self-Tuning
Decoupling the knowledge acquisition process into three perspectives is more effective than solely focusing on comprehension. The comparison between Self-Tuning w/o Review and Self-Tuning w/ Read demonstrates that presenting the test document text from three distinct perspectives contributes more to knowledge memorization (1.26% vs. 1.46% on PPL), extraction (23.68% vs. 17.65% on EM), and reasoning (50.40% vs. 39.37% on accuracy) than presenting the test document text with all constructed tasks as a whole.
@@ 5.7 Fine-grained Comparison
Results. As shown in Table 7, we observe that Self-Tuning consistently outperforms PIT in the overall evaluation and the fine-grained evaluations related to different QA types. These findings underscore the importance of equipping the model with the ability to systematically acquire new knowledge. Moreover, we present a qualitative comparison between the answers generated by PIT and Self-Tuning in Appendix D. Additionally, to provide insights for future work, we conduct a detailed error analysis on the types of factual errors that remain challenging after implementing Self-Tuning in Appendix E.
@@ Appendix B Evaluation Results of Additional Compared Methods
Setup. For a thorough assessment, we also examine the efficiency of our suggested Self-Tuning by contrasting it with four other notable methods, namely standard instruction-tuning without forgetting, QA adaptation, PIT, and mixed training, as displayed in Table 8.
Results. We present the evaluation results in Table A. Our Self-Tuning consistently shows superior performance, for example, it increases the EM by 11% on the knowledge extraction task.
In addition, Self-Tuning significantly outperforms QA adaptation in knowledge retention and extraction tasks, emphasizing the need to equip the model with the capability to systematically assimilate knowledge from raw documents instead of merely providing it with question-answering skills. Surprisingly, QA adaptation achieves high accuracy on the reasoning task. This could be attributed to an LLM’s ability to develop impressive reasoning skills during its pre-training phase, and a model with robust reasoning capabilities might excel in unfamiliar knowledge-intensive tasks Ovadia et al. (2024). Consequently, we regard knowledge extraction as the primary evaluation dimension, in line with previous studies Ovadia et al. (2024); Jiang et al. (2024b).
@@ Appendix E Error Analysis
| Paraphrase | 10.29% | <Siti Atiqoh Supriyanti - Wikipedia> Siti Atiqoh Supriyanti (born 25 November 1971) is a former civil servant and journalist who served as the First Lady of Central Java from 2013 to 2023. She is the wife of Ganjar Pranowo, a politician who is a member of Indonesian Indonesian Democratic Party of Struggle and Governor of Central Java. | When did Siti Atiqoh Supriyanti serve as the First Lady of Central Java? | 2013 to 2023. | From 2013 to 2023. |
To gain insights into potential enhancements for Self-Tuning, we outline four common errors that persist as challenges after implementing Self-Tuning, and offer an in-depth analysis in Table 12.
@@ Appendix I Detailed Templates used in the Self-Teaching Strategy
It is well-known that knowledge acquisition is often accompanied by catastrophic forgetting Allen-Zhu and Li (2023); Wang et al. (2023). Therefore, we also provide the knowledge retention performance for a comprehensive investigation. Specifically, we verify the knowledge extraction performance on world knowledge using natural questions (NQ) Kwiatkowski et al. (2019) (i.e., NQ-open Min et al. (2021) in the closed-book setting) and report EM and F1 scores. We report the five-shot evaluation results using the first five QA pairs in the dev sets as prompts. we assess the reasoning capability on Commonsense knowledge using CommonsenseQA (CSQA) Talmor et al. (2019), employing accuracy to assess the correctness of the selected option, calculated by comparing the generated option against the gold option using EM. We present the five-shot performance on the dev sets, as the test set does not contain golden annotations, and use the first five multi-choice QA pairs in the training set as prompts.
@@
QA adaptation first trains on and for 1 epoch, then on for 2 epochs.

### v3

@@ Abstract
Large language models (LLMs) often struggle to provide up-to-date information due to their one-time training and the constantly evolving nature of the world. To keep LLMs current, existing approaches typically involve continued pre-training on new documents. However, they frequently face difficulties in extracting stored knowledge. Motivated by the remarkable success of the Feynman Technique in efficient human learning, we introduce Self-Tuning, a learning framework aimed at improving an LLM’s ability to effectively acquire new knowledge from raw documents through self-teaching. Specifically, we develop a Self-Teaching strategy that augments the documents with a set of knowledge-intensive tasks created in a self-supervised manner, focusing on three crucial aspects: memorization, comprehension, and self-reflection. In addition, we introduce three Wiki-Newpages-2023-QA datasets to facilitate an in-depth analysis of an LLM’s knowledge acquisition ability concerning memorization, extraction, and reasoning. Extensive experimental results on Llama2 family models reveal that Self-Tuning consistently exhibits superior performance across all knowledge acquisition tasks and excels in preserving previous knowledge.††∗Work done during the internship at Tencent AI Lab. 111Our source code and the curated new datasets can be accessed publicly at https://github.com/zhangxy-2019/Effective-Knowledge-Injection.
@@ 1 Introduction
In this paper, we focus on injecting new knowledge into the parameters of LLMs. As depicted in the upper part of Figure 1, a standard approach involves continued pre-training (A) on a raw corpus (here, test doc) containing new information Jang et al. (2022). However, it struggles to extract the embedded knowledge, potentially due to the impaired question-answering (QA) capability Allen-Zhu and Li (2023); Cheng et al. (2024). Despite the assistance of subsequent instruction-tuning (B) Wei et al. (2022); Ouyang et al. (2022a) on QA data, the knowledge retrieved from the LLMs remains notably constrained Jiang et al. (2024b). Recently, Jiang et al. (2024b) suggests fine-tuning on a mix of QA data and related documents before continuing pre-training, with the aim of teaching the model how to access knowledge from documents and answer questions. Although this method greatly outperforms standard approaches, our initial results suggest that its effectiveness in knowledge extraction remains limited.
@@
We introduce three Wiki-Newpages-2023-QA datasets to enable a comprehensive analysis of an LLM’s knowledge acquisition ability w.r.t., memorization, extraction, and reasoning.
@@ 4.1 Overview
Our objective is to equip an LLM , parameterized by , with the ability to learn how to derive knowledge from raw documents. This is achieved by training the model using a combination of training document dataset and associated training QA dataset , as depicted in the upper left part of Figure 2. To enhance effective knowledge absorption, we present along with a series of knowledge-intensive tasks (a.k.a. self-teaching tasks) that are related to their content for Self-Teaching (in the lower part of Figure 2). These tasks are generated in a self-supervised manner based on the contents of , using the proposed Self-Teaching learning approach (Section 4.2). The multi-task training objective is presented as follows:
@@
Our aim is to train the model to apply the learned strategy for spontaneously extracting new knowledge from unseen documents (i.e., the test document dataset ). In addition to training on , we include , allowing the model to review and refine its question-answering ability. This approach enhances knowledge extraction from . The training objective is as follows:
@@
Our goal is to ensure that the model thoroughly absorbs the new knowledge from unseen documents by conducting follow-up training on the dataset. The objective for this stage is as follows:
@@ 5.3 Training Dynamics
We analyze the training dynamics of Self-Tuning during continued pre-training (beginning from Stage 2 in Figure 2) on the test documents by varying the number of training epochs for two main reasons: to eliminate the possibility that the exceptional performance of Self-Tuning in enhancing knowledge acquisition is merely a result of early fitting on the test documents, and to conduct an in-depth assessment of its long-term knowledge retention capability. Furthermore, we integrate the results of PIT and continued pre-training to offer a well-rounded evaluation.
@@ 5.5 Variants of Self-Tuning
Decoupling the knowledge acquisition process into three perspectives is more effective than solely focusing on comprehension. The comparison between Self-Tuning w/o Review and Self-Tuning w/ Read. demonstrates that presenting the test document text from three distinct perspectives contributes more to knowledge memorization (1.26% vs. 1.46% on PPL), extraction (23.68% vs. 17.65% on EM), and reasoning (50.40% vs. 39.37% on accuracy) than presenting the test document text with all constructed tasks as a whole.
@@ 5.7 Fine-grained Comparison
Results. As shown in Table 7, we observe that Self-Tuning consistently outperforms PIT in the overall evaluation and the fine-grained evaluations related to different QA types. These findings underscore the importance of equipping the model with the ability to systematically acquire new knowledge. Furthermore, we present a qualitative comparison between the answers generated by PIT and Self-Tuning in Appendix D. To gain insights into potential enhancements for Self-Tuning, we also conduct a detailed error analysis on the types of factual errors that remain challenging after implementing Self-Tuning in Appendix E.
@@ Appendix B Evaluation Results of Additional Compared Methods
For a thorough assessment, we also examine the efficiency of our suggested Self-Tuning by contrasting it with three other notable methods, namely standard instruction-tuning without forgetting, PIT++, and mixed training, as displayed in Table 8. We present the evaluation results in Table A. Our Self-Tuning consistently shows superior performance, for example, it increases the EM by 11% on the knowledge extraction task.
@@ Appendix E Error Analysis
| Paraphrase | 10.29% | <Lil Tay - Wikipedia> Tay Tian (born July 29, 2009), known professionally as Lil Tay, is an American-born Canadian internet personality and singer. In 2018, she gained prominence online for a period of three months, proclaiming herself to be the “youngest flexer of the century”. During her brief career, she posted rap videos on YouTube and Instagram which garnered tens of millions of views. Her career ended in mid-2018, after her father applied to the superior court of Canada for full custody and control of her career. According to court documents, he was abusive and largely an absentee. | What is Lil Tay’s nationality? | American-born Canadian. | Canadian-American. |
In order to gain insights into potential enhancements for Self-Tuning, we outline four common errors that persist as challenges after implementing Self-Tuning. We offer an in-depth analysis of these errors in Table 12, using EM as the evaluation metric.
@@ Appendix I Detailed Templates used in the Self-Teaching Strategy
It is well-known that knowledge acquisition is often accompanied by catastrophic forgetting Allen-Zhu and Li (2023); Wang et al. (2023). Therefore, we also provide the knowledge retention performance for a comprehensive investigation. Specifically, we verify the knowledge extraction performance on world knowledge using natural questions (NQ) Kwiatkowski et al. (2019) (i.e., NQ-open Min et al. (2021) in the closed-book setting) and report EM and F1 scores. We report the five-shot evaluation results using the first five QA pairs in the dev sets as prompts. we assess the reasoning capability on Commonsense knowledge using CommonsenseQA (CSQA) Talmor et al. (2019), employing accuracy to assess the correctness of the selected option, calculated by comparing the generated option against the gold option using EM. We present the five-shot performance on the dev sets, as the test set does not contain golden annotations, and use the first five multi-choice QA pairs in the training set as prompts. We use these two datasets because they were curated before the cut-off time of Llama2 family models (i.e., year 2022), making it likely that the models have obtained relevant knowledge in these datasets during the pre-training stage, as evidenced by Touvron et al. (2023).